# A Statistical Framework of Watermarks for Large Language Models: Pivot, Detection Efficiency and Optimal Rules

### v3

@@ B.3 Analysis of Different Score Functions
To visualize the differences among the tests induced by these $h$’s, we examine a two-sample two-token case (i.e., $n=2$ and $|{{\mathcal{W}}}|=2$). The left panel in Figure 8 depicts the $\alpha$-level rejection region for different $h$’s. The rejection region defined by $h_{\mathrm{llr}},h_{\mathrm{log}},h_{\mathrm{ind},\delta}$ exhibits a more concentrated pattern around the corner $(1,1)$, in contrast to the elongated and narrow shape defined by $h_{\mathrm{ars}}$. Furthermore, the rejection regions of $h_{\mathrm{llr}}$ and $h_{\mathrm{log}}$ share almost the same shape.
@@
$h_{\mathrm{llr}}(y)=\log f_{\bm{P}}(y)$ $-D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}})$ $\int_{0}^{1}f_{\bm{P}}^{1-\theta}(y){\mathrm{d}}y$ $D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}})$
$h_{\mathrm{ars}}(y)=-\log(1-y)$ 1 $\sum_{{w}\in{{\mathcal{W}}}}B(1/P_{{w}},\theta+1)$ $-\inf\limits_{\theta\geq 0}\left[\theta+\log\phi_{\bm{P},h_{\mathrm{ars}}}(\theta)\right]$
$h_{\mathrm{log}}(y)=\log y$ $-1$ $\sum_{{w}\in{{\mathcal{W}}}}\frac{P_{{w}}}{1-P_{{w}}\theta}$ $\sup\limits_{\theta\in[0,\frac{1}{P_{(1)}})}[\theta-\log\phi_{\bm{P},h_{\mathrm{log}}}(\theta)]$
@@
  • For $h_{\mathrm{llr}}(y)=\log f_{\bm{P}}(y)$, $-\operatorname{\mathbb{E}}_{0}h_{\mathrm{llr}}(Y)=R_{\bm{P}}(h_{\mathrm{llr}})=D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}})$.
  • For $h_{\mathrm{ars}}(y)=-\log(1-y)$, $\operatorname{\mathbb{E}}_{0}h_{\mathrm{ars}}(Y)=1$ and $\phi_{\bm{P},h_{\mathrm{ars}}}(\theta)=\sum_{{w}\in{{\mathcal{W}}}}B(1/P_{{w}},\theta+1)$ where $B(a,b)=\int_{0}^{1}y^{a-1}(1-y)^{b-1}{\mathrm{d}}y$ is the beta function. $\phi_{\bm{P},h_{\mathrm{ars}}}(\theta)$ is well-defined on $[0,\infty)$ and is strictly decreasing with $\phi_{\bm{P},h_{\mathrm{ars}}}(0)=\infty$ and $\phi_{\bm{P}}(h_{\mathrm{ars}})(\infty)=0$. Hence, $R_{\bm{P}}(h_{\mathrm{ars}})=-\theta_{1}^{\star}-\log\phi_{h_{\mathrm{ars}}}(\theta_{1}^{\star})$ where $\theta_{1}^{\star}$ is the unique solution of the equation $\phi_{\bm{P},h_{\mathrm{ars}}}(\theta)+\phi_{h_{\mathrm{ars}}}^{\prime}(\theta)=0$.
  • For $h_{\mathrm{log}}(y)=\log y$, $\operatorname{\mathbb{E}}_{0}h_{\mathrm{log}}(Y)=-1$ and $\phi_{\bm{P},h_{\mathrm{log}}}(\theta)=\sum_{{w}\in{{\mathcal{W}}}}\frac{P_{{w}}}{1-P_{{w}}\theta}$ which is well-defined on $[0,1/P_{(1)})$. Hence, $R_{\bm{P}}(h_{\mathrm{log}})=\theta_{2}^{\star}-\log\phi_{\bm{P},h_{\mathrm{log}}}(\theta_{2}^{\star})$ where $\theta_{2}^{\star}$ is the unique solution of the equation $\phi_{\bm{P},h_{\mathrm{log}}}(\theta)=\phi_{\bm{P},h_{\mathrm{log}}}^{\prime}(\theta)$.
  • For $h_{\mathrm{ind},\delta}(y)={\bf 1}_{\{y\geq\delta\}}$ for some $\delta\in(0,1)$, $\operatorname{\mathbb{E}}_{0}h_{\mathrm{ind},\delta}(Y)=1-\delta$ and $\phi_{\bm{P},h_{\mathrm{ind},\delta}}(\theta)=F(\delta)+{\mathrm{e}}^{-\theta}(1-F(\delta))$ which is well-defined on $[0,\infty)$. Hence, $R_{P}(h_{\mathrm{ind},\delta})=\delta\log\frac{\delta}{F(\delta)}+(1-\delta)\log\frac{1-\delta}{1-F(\delta)}$.
@@
  • Suppose that we already show the statement holds for $m\leq\lfloor\frac{1}{1-\Delta}\rfloor$, we now need to show that it also holds for $m-1$. For notation simplicity, we let ${\mathcal{P}}_{m}\subset{\mathcal{P}}_{\Delta}$ denote the set of NTP distributions which having exact $m$ coordinates equal to $1-\Delta$. Let $\bm{P}\in{\mathcal{P}}_{m-1}$ be the NTP distribution under consideration. Given its coordinates are decreasingly ranked, we must have $1-\Delta>P_{m}\geq P_{m+1}\geq\cdots\geq P_{|{{\mathcal{W}}}|}$.
@@
      In this case, though $\bm{P}_{1}$ and $\bm{P}_{2}$ are still in ${\mathcal{P}}_{m-1}$, their support decreases by one when compared with $\bm{P}$. We then discuss the remaining coordinates of $\bm{P}_{1}$ and $\bm{P}_{2}$. Take $\bm{P}_{1}$ for example. We want to determine whether the sum of the largest two non-$(1-\Delta)$ probabilities in $\bm{P}_{1}$ is larger than $1-\Delta$ or not. If this is true, we arrive at Case (a) again but with a new instance, namely $\bm{P}_{1}$, which has already been addressed. If this is not true, we can further decrease the support of $\bm{P}_{1}$ by treating it as a new instance in Case (b) and combining these two largest non-$(1-\Delta)$ coordinates. Because $\bm{P}_{1}$ has a finite number of non-$(1-\Delta)$ coordinates, there is only a finite number of cases where we have to deal with Case (b). As a result, by tracing the way we decompose $\bm{P}_{1}$, we always have that $\bm{P}_{1}\in\mathrm{Conv}(\Pi(\bm{P}_{\Delta}^{\star}))$ and similarly $\bm{P}_{2}\in\mathrm{Conv}(\Pi(\bm{P}_{\Delta}^{\star}))$.
@@ B.5 Results for a General Probability Set
In the following, we consider a new, general distribution class ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$. For two real numbers $10\leq\Delta_{2}<\Delta_{1}<$ 1, we define the following probability class:
@@
For $10\leq\Delta_{2}<\Delta_{1}<$ 1, the set of extreme points of ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$ is
@@
There exist two positive constants $10.001<\Delta_{1}^{{\mathrm{gumb}}}\leq\Delta_{2}^{{\mathrm{gumb}}}<$ 1 such that
@@
[…]
@@
$R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{ars}})$ $=-\inf\limits_{\theta\geq 0}\left[\theta+\log\phi_{\bm{P}^{\star}_{\Delta},h_{\mathrm{ars}}}(\theta)\right]:=R_{h_{\mathrm{ars}}}(\Delta),$
$R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{log}})$ $=\sup\limits_{\theta\in[0,\frac{1}{1-\Delta})}[\theta-\log\phi_{\bm{P}^{\star}_{\Delta},h_{\mathrm{log}}}(\theta)]:=R_{h_{\mathrm{log}}}(\Delta),$
$R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{ind},\delta})$ $=\delta\log\frac{\delta}{F_{1,\bm{P}_{\Delta}^{\star}}(\delta)}+(1-\delta)\log\frac{1-\delta}{1-F_{1,\bm{P}_{\Delta}^{\star}}(\delta)}:=R_{h_{\mathrm{ind},\delta}}(\Delta),$
@@ B.7 Analysis of Efficiency Gap
$1-\min\left\{\frac{P_{(1)}}{1-P_{(1)}}\log\frac{1}{P_{(1)}},\frac{|\mathrm{supp}(\bm{P})|-\frac{1}{P_{(1)}}}{|\mathrm{supp}(\bm{P})|-1-\log|\mathrm{supp}(\bm{P})|}\right\}\leq\frac{R_{\bm{P}}(h_{\mathrm{log}})}{R_{\bm{P}}(h_{\mathrm{llr}})}\leq 1$
@@
The last equality uses the condition $P_{(1)}<1$ so that the supreme is obtained within $[0,1/P_{(1)})$. On the other hand, note that $\lim\limits_{p\to 0}r^{1/p-1}=0$ for any $r\in(0,1)$ and thus
@@ B.8 ${\mathcal{P}}_{\Delta}$-Efficiency for the Baby Watermark
Note that for $y<-1$, $(y+1)/2<0$ and $(1-y)/2>1$ and for $y>1$, $(y+1)/2>1$ and $(1-y)/2<0$. Thus, ${\mathbb{P}}(Y\leq y)=0$ for $y<-1$ and ${\mathbb{P}}(Y\leq y)=1$ for $y>1$. For $-1\leq y\leq 1$, it suffices to calculate ${\mathbb{P}}(\zeta\leq a\mid{w}=1)$ and ${\mathbb{P}}(\zeta\leq a\mid{w}=0)$ for any $-1\leq a\leq 1$. By the Bayes rule,
@@
for $-1\leq y\leq 1$. If $P_{0}\geq 1/2$, (35) holds with $P_{0}$ being replaced by $P_{1}$. We then complete the proof. ∎
@@ C.1 Distribution Characterization
where $a_{\pi,{w}}=\sum_{j=1}^{{w}}P_{t,\pi(j)}$ is the sum of the first ${w}$ probabilities of $\bm{P}_{t}$ under the permutation $\pi$.
@@
where $a_{\pi^{-1},{w}}=\sum_{j\in{{\mathcal{W}}}}p_{t,\pi^{-1}(j)}$ is the endpoint of $S_{t}(\pi^{-1}({w}),\pi)$ by definition. The last equation uses $\pi^{-1}\in\Pi$ is equivalent to $\pi\in\Pi$.
@@ C.2 Key Lemma for Asymptotic Analysis
Let $\mathcal{J}$ collect any $1$-Lipschitz-continuous function on $[0,1]^{2}$. For any $J\in\mathcal{J}$, $J(U_{t},\eta(\pi_{t}({w}_{t})))$ is a bounded random variable where $\eta({w})=\frac{{w}-1}{|{{\mathcal{W}}}|-1}$. Let $\bm{P}_{t}$ be the NTP distribution that ${w}_{t}$ follows. We introduce a new belief class, denoted by $\mathcal{Q}_{\Delta}$ with the following definition
@@
It suffices to show (39) holds for any indicator functions, i.e., $J(x,y)={\bf 1}_{\{x\leq r_{1},y\leq r_{2}\}}$ for a given pair $(r_{1},r_{2})\in[0,1]^{2}$. For the indicator function $x\mapsto{\bf 1}_{\{x\leq r\}}$, we introduce its Lipschitz variant $f_{{\varepsilon},r}(x)$ for a small number ${\varepsilon}>0$:
$f_{{\varepsilon},r}(x):=\begin{cases}1&~{}\text{if}~{}0\leq x\leq r,\\ 1-\frac{x-r}{{\varepsilon}}&~{}\text{if}~{}r\leq x\leq r+{\varepsilon},\\ 0&~{}\text{if}~{}r+{\varepsilon}\leq x\leq 1.\\ \end{cases}$
@@
This above proof yields an insightful result. Employing $J(x,y)={\bf 1}_{\{x\leq r_{1},y\leq r_{2}\}}$ within (39) reveals the asymptotic CDF of $(U_{t},\eta(\pi_{t}({w}_{t})))$ to be free of discontinuities across the square $[0,1]^{2}$. This continuity is a direct consequence of the right-hand side of (39) (substituting $J(x,y)={\bf 1}_{\{x\leq r_{1},y\leq r_{2}\}}$) being continuous in $(r_{1},r_{2})$.
@@
where $a_{\pi,{w}}=\sum_{j=1}^{{w}}P_{t,\pi(j)}$ is the sum of the first ${w}$ probabilities of $\bm{P}_{t}$ under the permutation $\pi$. Hence,
@@
Since we fix $t\in[n]$, we drop the dependence of $\bm{P}_{t}$ on $t$ for simplicity. We let $P_{(1)}=\max_{{w}\in{{\mathcal{W}}}}P_{{w}}$ denote the largest probability in $P$ so that $P_{(1)}=1-\Delta$. Now, we discuss the position, denoted by $\ell$, that is mapped to 1 according to the permutation $\pi_{t}$, i.e., $\pi_{t}(\ell)=1$ and divide the summation $\sum_{\pi\in\Pi}$ into three parts. We use $\Pi_{\ell}$ to represent the set of permutations that map $\ell$ to 1. Then $\Pi=\bigcup_{\ell\in{{\mathcal{W}}}}\Pi_{\ell}$ and thus
@@
$\quad+\frac{1}{|{{\mathcal{W}}}|!}\sum_{{w}\in{{\mathcal{W}}}}\sum_{\pi\in\Pi_{k}}\int_{a_{\pi,{w}-1}}^{a_{\pi,{w}}}J(r,\eta({w})){\mathrm{d}}r$
@@
Note that given the condition $\bar{\pi}(\ell)=1$, $\bar{\pi}$ is still a random permutation on the resting $|{{\mathcal{W}}}|-1$ numbers which permutes ${{\mathcal{W}}}-\{\ell\}$ to ${{\mathcal{W}}}-\{1\}$. To that end, we will use the concentration inequality for randomly permuted sums developed by [4].
@@
Let $b_{j,l}\geq 0$ be a collection of non-negative numbers and $\bar{\pi}$ be a random uniform permutation in ${{\mathcal{W}}}$, i.e., $\bar{\pi}\sim{U}({{\mathcal{W}}})$. Let $Z=\sum_{j\in{{\mathcal{W}}}}b_{j,\bar{\pi}(j)}$. Then for any $x>0$,
@@
However, whether the finite-vocabulary alternative CDF, i.e., $F_{1,\bm{P}_{t}}^{{\mathrm{dif}}}$, converges to its limit CDF depends on whether the condition $\log|{{\mathcal{W}}}|\cdot P_{t,(1)}=o(1)$ holds. For example, if $\bm{P}_{t}\sim\mathrm{Pow}(2,1)$, then $\log|{{\mathcal{W}}}|\cdot P_{t,(1)}=\Omega(1)$, and consequently, the finite-vocabulary alternative CDF does not converge to the limit (see the purple dotted curves in Figure 9). In contrast, if $\bm{P}_{t}$ follows $\mathrm{Pow}(a,b)$ with $a<1$, then $\log|{{\mathcal{W}}}|\cdot P_{t,(1)}=o(1)$ is satisfied, so the finite-vocabulary CDF aligns well with its limit even if $|{{\mathcal{W}}}|=100$.
@@
$\lim_{|{{\mathcal{W}}}|\to\infty}\left|\sup_{\Delta^{\prime}\geq\Delta}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\phi_{\bm{P},h}(1)-\sup_{\Delta^{\prime}\geq\Delta}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\lim_{|{{\mathcal{W}}}|\to\infty}\phi_{\bm{P},h}(1)\right|=0.$
Note that by the uniform convergence (46), we have
$\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\lim_{|{{\mathcal{W}}}|\to\infty}\phi_{\bm{P},h}(1)=\lim_{|{{\mathcal{W}}}|\to\infty}\phi_{\bm{P},h}(1)=\lim_{|{{\mathcal{W}}}|\to\infty}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\phi_{\bm{P},h}(1).$ (47)
The last two equations imply that
@@
[…]
@@
[…]

### v4

@@ B.3 Analysis of Different Score Functions
To visualize the differences among the tests induced by these $h$’s, we examine a two-sample two-token case (that is, $n=2$ and $|{{\mathcal{W}}}|=2$). The left panel in Figure 8 depicts the $\alpha$-level rejection region for different $h$’s. The rejection region defined by $h_{\mathrm{llr}},h_{\mathrm{log}},h_{\mathrm{ind},\delta}$ exhibits a more concentrated pattern around the corner $(1,1)$, in contrast to the elongated and narrow shape defined by $h_{\mathrm{ars}}$. Furthermore, the rejection regions of $h_{\mathrm{llr}}$ and $h_{\mathrm{log}}$ share almost the same shape.
@@
$h_{\mathrm{llr}}(y)=\log f_{\bm{P}}(y)$ $D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}})$ $\int_{0}^{1}f_{\bm{P}}^{1-\theta}(y){\mathrm{d}}y$ $D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}})$
$h_{\mathrm{ars}}(y)=-\log(1-y)$ 1 $\sum_{{w}\in{{\mathcal{W}}}}B(1/P_{{w}},\theta+1)$ $\inf\limits_{\theta\geq 0}\left[\theta+\log\phi_{\bm{P},h_{\mathrm{ars}}}(\theta)\right]$
$h_{\mathrm{log}}(y)=\log y$ 1-1 $\sum_{{w}\in{{\mathcal{W}}}}\frac{P_{{w}}}{1-P_{{w}}\theta}$ $\sup\limits_{\theta\in[0,\frac{1}{P_{(1)}})}[\theta-\log\phi_{\bm{P},h_{\mathrm{log}}}(\theta)]$
@@
  • For $h_{\mathrm{llr}}(y)=\log f_{\bm{P}}(y)$, $\operatorname{\mathbb{E}}_{0}h_{\mathrm{llr}}(Y)=R_{\bm{P}}(h_{\mathrm{llr}})=D_{\mathrm{KL}}(\mu_{0},\mu_{1,\bm{P}})$.
  • For $h_{\mathrm{ars}}(y)=-\log(1-y)$, $\operatorname{\mathbb{E}}_{0}h_{\mathrm{ars}}(Y)=1$ and $\phi_{\bm{P},h_{\mathrm{ars}}}(\theta)=\sum_{{w}\in{{\mathcal{W}}}}B(1/P_{{w}},\theta+1)$ where $B(a,b)=\int_{0}^{1}y^{a-1}(1-y)^{b-1}{\mathrm{d}}y$ is the beta function. $\phi_{\bm{P},h_{\mathrm{ars}}}(\theta)$ is well-defined on $0,\infty)$ and is strictly decreasing with $\phi_{\bm{P},h_{\mathrm{ars}}}(0)=\infty$ and $\phi_{\bm{P}}(h_{\mathrm{ars}})(\infty)=0$. Hence, $R_{\bm{P}}(h_{\mathrm{ars}})=-\theta_{1}^{\star}-\log\phi_{h_{\mathrm{ars}}}(\theta_{1}^{\star})$ where $\theta_{1}^{\star}$ is the unique solution of the equation $\phi_{\bm{P},h_{\mathrm{ars}}}(\theta)+\phi_{h_{\mathrm{ars}}}^{\prime}(\theta)=0$.
  • For $h_{\mathrm{log}}(y)=\log y$, $\operatorname{\mathbb{E}}_{0}h_{\mathrm{log}}(Y)=-1$ and $\phi_{\bm{P},h_{\mathrm{log}}}(\theta)=\sum_{{w}\in{{\mathcal{W}}}}\frac{P_{{w}}}{1-P_{{w}}\theta}$ which is well-defined on $0,1/P_{(1)})$. Hence, $R_{\bm{P}}(h_{\mathrm{log}})=\theta_{2}^{\star}-\log\phi_{\bm{P},h_{\mathrm{log}}}(\theta_{2}^{\star})$ where $\theta_{2}^{\star}$ is the unique solution of the equation $\phi_{\bm{P},h_{\mathrm{log}}}(\theta)=\phi_{\bm{P},h_{\mathrm{log}}}^{\prime}(\theta)$.
  • For $h_{\mathrm{ind},\delta}(y)={\bf 1}_{\{y\geq\delta\}}$ for some $\delta\in(0,1)$, $\operatorname{\mathbb{E}}_{0}h_{\mathrm{ind},\delta}(Y)=1-\delta$ and $\phi_{\bm{P},h_{\mathrm{ind},\delta}}(\theta)=F(\delta)+{\mathrm{e}}^{-\theta}(1-F(\delta))$ which is well-defined on $0,\infty)$. Hence, $R_{P}(h_{\mathrm{ind},\delta})=\delta\log\frac{\delta}{F(\delta)}+(1-\delta)\log\frac{1-\delta}{1-F(\delta)}$.
@@
This lemma can be proved by directly applying Lemma B.4 using the concept of majorization. Below, we provide a constructive proof as an alternative.
@@
  • Suppose that we already show the statement holds for $m\leq\lfloor\frac{1}{1-\Delta}\rfloor$, we now need to show that it also holds for m1m-1. For notation simplicity, we let ${\mathcal{P}}_{m}\subset{\mathcal{P}}_{\Delta}$ denote the set of NTP distributions which having exact $m$ coordinates equal to $1-\Delta$. Let $\bm{P}\in{\mathcal{P}}_{m-1}$ be the NTP distribution under consideration. Given its coordinates are decreasingly ranked, we must have $1-\Delta>P_{m}\geq P_{m+1}\geq\cdots\geq P_{|{{\mathcal{W}}}|}$.
@@
[…]
@@ B.5 Results for a General Probability Set
In the following, we consider a new, general distribution class ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$. For two real numbers $0\leq\Delta_{2}<\Delta_{1}<1$, we define the following probability class:
@@
For $0\leq\Delta_{2}<\Delta_{1}<1$, the set of extreme points of ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$ is
@@
There exist two positive constants $0.001<\Delta_{1}^{{\mathrm{gumb}}}\leq\Delta_{2}^{{\mathrm{gumb}}}<1$ such that
@@
  • From Figure 4, we observe that $\Delta_{1}^{{\mathrm{gumb}}}=\Delta_{2}^{{\mathrm{gumb}}}\approx 0.17756080525215662$.
@@
$R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{ars}})$ $=-\inf\limits_{\theta\geq 0}\left[\theta+\log\phi_{\bm{P}^{\star}_{\Delta},h_{\mathrm{ars}}}(\theta)\right]=R_{h_{\mathrm{ars}}}(\Delta),$
$R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{log}})$ $=\sup\limits_{\theta\in[0,\frac{1}{1-\Delta})}[\theta-\log\phi_{\bm{P}^{\star}_{\Delta},h_{\mathrm{log}}}(\theta)]=R_{h_{\mathrm{log}}}(\Delta),$
$R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{ind},\delta})$ $=\delta\log\frac{\delta}{F_{1,\bm{P}_{\Delta}^{\star}}(\delta)}+(1-\delta)\log\frac{1-\delta}{1-F_{1,\bm{P}_{\Delta}^{\star}}(\delta)}=R_{h_{\mathrm{ind},\delta}}(\Delta),$
@@ B.7 Analysis of Efficiency Gap
$1-\min\left\{\frac{P_{(1)}}{1-P_{(1)}}\log\frac{1}{P_{(1)}},\frac{|\mathrm{supp}(\bm{P})|-\frac{1}{P_{(1)}}}{|\mathrm{supp}(\bm{P})|-1-\log|\mathrm{supp}(\bm{P})|}\right\}\leq\frac{R_{\bm{P}}(h_{\mathrm{log}})}{R_{\bm{P}}(h_{\mathrm{llr}})}\leq 1.$
@@
The last equality uses the condition $P_{(1)}<1$ so that the supreme is obtained within $0,1/P_{(1)})$. On the other hand, note that $\lim\limits_{p\to 0}r^{1/p-1}=0$ for any $r\in(0,1)$ and thus
@@ B.8 ${\mathcal{P}}_{\Delta}$-Efficiency for the Baby Watermark
Note that for y<1y<-1, $(y+1)/2<0$ and (1y)/2>1(1-y)/2>1 and for $y>1$, $(y+1)/2>1$ and (1y)/2<0(1-y)/2<0. Thus, ${\mathbb{P}}(Y\leq y)=0$ for y<1y<-1 and ${\mathbb{P}}(Y\leq y)=1$ for $y>1$. For $1\leq y\leq 1$, it suffices to calculate ${\mathbb{P}}(\zeta\leq a\mid{w}=1)$ and ${\mathbb{P}}(\zeta\leq a\mid{w}=0)$ for any $1\leq a\leq 1$. By the Bayes rule,
@@
for $1\leq y\leq 1$. If $P_{0}\geq 1/2$, (35) holds with $P_{0}$ being replaced by $P_{1}$. We then complete the proof. ∎
@@ C.1 Distribution Characterization
where $a_{\pi,{w}}=\sum_{j=1}^{{w}}P_{t,\pi(j)}$ is the sum of the first w{w} probabilities of $\bm{P}_{t}$ under the permutation $\pi$.
@@
where $a_{\pi^{-1},{w}}=\sum_{j\leq{w}}P_{t,\pi^{-1}(j)}$ is the endpoint of $S_{t}(\pi^{-1}({w}),\pi)$ by definition. The last equation uses $\pi^{-1}\in\Pi$ is equivalent to $\pi\in\Pi$.
@@ C.2 Key Lemma for Asymptotic Analysis
Let $\mathcal{J}$ collect any $1$-Lipschitz-continuous function on $0,1]^{2}$. For any $J\in\mathcal{J}$, $J(U_{t},\eta(\pi_{t}({w}_{t})))$ is a bounded random variable where $\eta({w})=\frac{{w}-1}{|{{\mathcal{W}}}|-1}$. Let $\bm{P}_{t}$ be the NTP distribution that ${w}_{t}$ follows. We introduce a new belief class, denoted by $\mathcal{Q}_{\Delta}$ with the following definition
@@
It suffices to show (39) holds for any indicator functions, that is, $J(x,y)={\bf 1}_{\{x\leq r_{1},y\leq r_{2}\}}$ for a given pair $(r_{1},r_{2})\in[0,1]^{2}$. For the indicator function $x\mapsto{\bf 1}_{\{x\leq r\}}$, we introduce its Lipschitz variant $f_{{\varepsilon},r}(x)$ for a small number ${\varepsilon}>0$:
[…]
@@
This above proof yields an insightful result. Employing $J(x,y)={\bf 1}_{\{x\leq r_{1},y\leq r_{2}\}}$ within (39) reveals the asymptotic CDF of $(U_{t},\eta(\pi_{t}({w}_{t})))$ to be free of discontinuities across the square $0,1]^{2}$. This continuity is a direct consequence of the right-hand side of (39) (substituting $J(x,y)={\bf 1}_{\{x\leq r_{1},y\leq r_{2}\}}$) being continuous in $(r_{1},r_{2})$.
@@
where $a_{\pi,{w}}=\sum_{j=1}^{{w}}P_{t,\pi(j)}$ is the sum of the first w{w} probabilities of $\bm{P}_{t}$ under the permutation $\pi$. Hence,
@@
Since we fix $t\in[n]$, we drop the dependence of $\bm{P}_{t}$ on $t$ for simplicity. We let $P_{(1)}=\max_{{w}\in{{\mathcal{W}}}}P_{{w}}$ denote the largest probability in $P$ so that $P_{(1)}=1-\Delta$. Now, we discuss the position, denoted by $\ell$, that is mapped to 1 according to the permutation $\pi_{t}$, that is, $\pi_{t}(\ell)=1$ and divide the summation $\sum_{\pi\in\Pi}$ into three parts. We use $\Pi_{\ell}$ to represent the set of permutations that map $\ell$ to 1. Then $\Pi=\bigcup_{\ell\in{{\mathcal{W}}}}\Pi_{\ell}$ and thus
@@
$\quad+\frac{1}{|{{\mathcal{W}}}|!}\sum_{{w}\in{{\mathcal{W}}}}\sum_{\pi\in\Pi_{{w}}}\int_{a_{\pi,{w}-1}}^{a_{\pi,{w}}}J(r,\eta({w})){\mathrm{d}}r$
@@
Note that given the condition $\bar{\pi}(\ell)=1$, $\bar{\pi}$ is still a random permutation on the resting $|{{\mathcal{W}}}|-1$ numbers which permutes $\mathcal{W}$$\{\ell\}$ to ${{\mathcal{W}}}-\{1\}$. To that end, we will use the concentration inequality for randomly permuted sums developed by [4].
@@
Let $b_{j,l}\geq 0$ be a collection of non-negative numbers and $\bar{\pi}$ be a random uniform permutation in $\mathcal{W}$}}, that is, $\bar{\pi}\sim{U}({{\mathcal{W}}})$. Let $Z=\sum_{j\in{{\mathcal{W}}}}b_{j,\bar{\pi}(j)}$. Then for any $x>0$,
@@
However, whether the finite-vocabulary alternative CDF, that is, $F_{1,\bm{P}_{t}}^{{\mathrm{dif}}}$, converges to its limit CDF depends on whether the condition $\log|{{\mathcal{W}}}|\cdot P_{t,(1)}=o(1)$ holds. For example, if $\bm{P}_{t}\sim\mathrm{Pow}(2,1)$, then $\log|{{\mathcal{W}}}|\cdot P_{t,(1)}=\Omega(1)$, and consequently, the finite-vocabulary alternative CDF does not converge to the limit (see the purple dotted curves in Figure 9). In contrast, if $\bm{P}_{t}$ follows $\mathrm{Pow}(a,b)$ with $a<1$, then $\log|{{\mathcal{W}}}|\cdot P_{t,(1)}=o(1)$ is satisfied, so the finite-vocabulary CDF aligns well with its limit even if $|{{\mathcal{W}}}|=100$.
@@
$\lim_{|{{\mathcal{W}}}|\to\infty}$ $\left|\sup_{\Delta^{\prime}\geq\Delta}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\phi_{\bm{P},h}(1)-\sup_{\Delta^{\prime}\geq\Delta}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\lim_{|{{\mathcal{W}}}|\to\infty}\phi_{\bm{P},h}(1)\right|$
$\leq\lim_{|{{\mathcal{W}}}|\to\infty}\sup_{\Delta^{\prime}\geq\Delta}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\left|\phi_{\bm{P},h}(1)-\lim_{|{{\mathcal{W}}}|\to\infty}\phi_{\bm{P},h}(1)\right|=0.$
Therefore, it follows that
$\lim_{|{{\mathcal{W}}}|\to\infty}\sup_{\Delta^{\prime}\geq\Delta}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\phi_{\bm{P},h}(1)=\sup_{\Delta^{\prime}\geq\Delta}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\lim_{|{{\mathcal{W}}}|\to\infty}\phi_{\bm{P},h}(1).$ (47)
Note that by the uniform convergence (46), we have for any fixed $\Delta^{\prime}$,
$\lim_{|{{\mathcal{W}}}|\to\infty}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\left|\phi_{\bm{P},h}(1)-\lim_{|{{\mathcal{W}}}|\to\infty}\phi_{\bm{P},h}(1)\right|=0$
so that
$\lim_{|{{\mathcal{W}}}|\to\infty}\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\phi_{\bm{P},h}(1)=\sup_{\bm{P}\in\mathcal{Q}_{\Delta^{\prime}}}\lim_{|{{\mathcal{W}}}|\to\infty}\phi_{\bm{P},h}(1).$ (48)
The equations (47) and (48) imply that
@@
Given that the limit exists, the $\limsup$ also exists and equals the limit. ∎
@@
As implied in the proof of Lemma 4.4, we can replace $\limsup$ with $\lim$. By Theorem 4.1, we already have